OpenStudy (anonymous):

A student opens a mathematics book to two facing pages. The product of the page numbers is 1482. Find page numbers for first and second page?

OpenStudy (anonymous):

what does it mean by "facing pages"?

OpenStudy (anonymous):

I'm not sure that is the question they gave me and have no clue by what it means either

OpenStudy (dumbcow):

second = first +1 let x be the second page x*(x-1) = 1482 x^2 -x -1482 = 0 (x-39)(x+38) =0 x = 39 first = 38, second=39

OpenStudy (dumbcow):

it means they are consecutive pages

OpenStudy (anonymous):

two facing pages means consecutive?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

Thank you for thay advice I didnt know that and my material Ive been reading over never gave that. THANK YOU !!!

OpenStudy (dumbcow):

your welcome